Calculate the number of chloride ions in 200 cm cube of 0.5 mol magnesium chloride solution

1
Expert's answer
2022-04-11T17:05:11-0400

Solution:

Magnesium Chloride = MgCl2


MgCl2 → Mg2+ + 2Cl

According to stoichiometry:

1 mol of MgCl2 produces 2 mol of Cl

Thus, 0.5 mol of MgCl2 produces:

(0.5 mol MgCl2) × (2 mol Cl / 1 mol MgCl2) = 1 mol Cl


One mole of any substance contains 6.022×1023 particles.

Therefore,

Number of Clions = (1 mol) × (6.022×1023 ions / 1 mol) = 6.022×1023 ions

Number of Clions = 6.022×1023


Answer: 6.022×1023 of chloride ions